# SEA: Spectral Edge Attack

### v2

@@ Abstract
Graph-based machine learning algorithms occupy an important position in today’s AI landscape. The ability of graph topology to represent complex data structures is both the key strength of graph algorithms and a source of their vulnerability. In other words, attacking or perturbing a graph can severely degrade the performance of graph-based methods. For the attack methods, the greatest challenge is achieving strong attack effectiveness while remaining undetected. To address this problem, this paper proposes a new attack model that employs spectral adversarial robustness evaluation to quantitatively analyze the vulnerability of each edge in a graph under attack. By precisely targeting the weakest links, the proposed approach achieves the maximum attack impact with minimal perturbation. Experimental results demonstrate the effectiveness of the proposed method.
@@ 1 Introduction
Many important datasets are naturally represented as graphs, spanning online social and recommendation platforms (Facebook, Amazon), molecular interaction resources such as BioGRID, and citation or document-link networks like PubMed and Arxiv. Given the prevalence of such data, a number of widely used tasks and algorithms have been developed accordingly. From the task perspective, the most common problems include classification and clustering. From the algorithmic or modeling perspective, representative approaches include graph neural networks and spectral clustering.
The broad deployment of these tasks and algorithms also makes their security a critical concern. As stated in the ancient Chinese military treatise The Art of War, “know yourself and know your enemy, and you will never be defeated.” Without strong, well-designed attack methodologies, it is difficult to meaningfully evaluate defensive techniques or drive their improvement.
Motivated by this perspective, this paper proposes a new high-strength attack method. The key idea is to perform spectral adversarial robustness evaluation on the graph structure before launching the attack, thereby quantitatively estimating the vulnerability of each edge. This enables the attacker to pinpoint and target the weakest links, achieving maximal impact with only a very small number of perturbations to undermine the system while remaining difficult to detect.
@@ 4 Experiments
In this section, we use the method proposed in this paper to attack two types of dominant models in the field of graph-based machine learning: graph neural networks and spectral clustering. For the selected edges to be attacked, our algorithm randomly assigns their weights to an integer value within a specified range.To demonstrate the effectiveness of our method, we use random sampling to select the same number of edges from the graph as an attack baseline. We then perturb their weights by assigning random values drawn from the same range used in our method, and compare the resulting attack impact against ours.
4.1 Attacking GNNs
The results are shown in Table 1
Table 1: Baseline test accuracy and test accuracy drops under edge-weight randomization attacks (Spade-selected vs. Random-selected). Random results are averaged over 20 runs (mean $\pm$ std).
Dataset Model Baseline (Test) Drop (Ours) Drop (Random, mean $\pm$ std) Ours (wins/20)
Cora GAT 0.8826 0.0367 0.0138 $\pm$ 0.0083 20/20
Cora GraphSAGE 0.8936 0.0202 0.0114 $\pm$ 0.0074 18/20
PubMed GAT 0.8058 0.0124 0.0090 $\pm$ 0.0032 18/20
PubMed GraphSAGE 0.8489 0.0061 0.0038 $\pm$ 0.0018 18/20
4.2 Attacking Spectral Clustering
The results are shown in Table 2
Table 2: USPS data set: ACC drops (percentage points). Larger drop indicates a stronger attack.
change_ratio Baseline ACC Our Drop Random Drop (mean $\pm$ std) Ours (wins/20)
0.01 66.4229 +3.5492 $1.9542\pm 3.8065$ 20/20
0.02 66.4121 +2.1402 $1.2363\pm 3.5204$ 20/20
0.03 66.4229 +1.7853 $1.4390\pm 2.9963$ 20/20
0.04 66.4229 +3.4846 $0.6458\pm 0.4842$ 20/20
0.05 66.4121 +3.3125 $1.2266\pm 2.8007$ 20/20
@@ 5 Conclusion
This paper proposes a novel method for precisely attacking the graphs used in graph-based machine learning algorithms, based on spectral adversarial robustness evaluation. Experiments on representative graph algorithms from neural networks and statistical learning—namely graph neural networks (GNNs) and spectral clustering—demonstrate that the proposed method is highly effective.

### v3

@@ Abstract
Graph neural networks (GNNs) have been widely applied in a variety of domains. However, the very ability of graphs to represent complex data structures is both the key strength of GNNs and a major source of their vulnerability. Recent studies have shown that attacking GNNs by maliciously perturbing the underlying graph can severely degrade their performance. For attack methods, the central challenge is to maintain attack effectiveness while remaining difficult to detect. Most existing attacks require modifying the graph structure, such as adding or deleting edges, which is relatively easy to notice. To address this problem, this paper proposes a new attack model that employs spectral adversarial robustness evaluation to quantitatively analyze the vulnerability of each edge in a graph. By precisely targeting the weakest links, our method can achieve effective attacks without changing the connectivity pattern of edges in the graph, for example by subtly adjusting the weights of a small subset of the most vulnerable edges. We apply the proposed method to attack several classical graph neural network architectures, and experimental results show that our attack is highly effective.
@@ 1 Introduction
Many important datasets are naturally represented as graphs, including online social and recommendation platforms (e.g., Facebook, Amazon), molecular interaction databases such as BioGRID, and citation or document-link networks like PubMed and Arxiv. Given the prevalence of graph-structured data, a wide range of tasks and algorithms have been developed. Among these tasks, node classification is particularly common, with applications ranging from annotating proteins in biological interaction networks [6] and profiling users in e-commerce graphs [7] to categorizing scientific articles in citation networks [8]. From an algorithmic perspective, graph neural networks (GNNs) are among the most widely used approaches [10].
GNNs excel at exploiting relational information by propagating and aggregating signals over the underlying graph, thereby learning task-relevant node representations that explicitly incorporate graph topology. In contrast, conventional neural networks typically operate on independent feature vectors and therefore cannot directly benefit from the structural dependencies encoded by edges. However, this very strength of GNNs is also the root of their vulnerability to adversarial manipulation [12, 13]. Prior work has demonstrated that even a small number of graph topology changes—such as inserting, deleting, or rewiring a few edges—can lead to substantial performance degradation [11, 14].
Given the widespread use of GNNs and their susceptibility to adversarial perturbations of the graph structure [11], their security has become a critical concern. To systematically test and analyze their robustness, many adversarial attack methods against GNNs have been proposed. Most existing adversarial attack methods on graphs focus on reducing the predictive accuracy of GNN models by subtly inserting or removing edges, often under constraints designed to preserve properties such as the node degree distribution [1]. Broadly, these attacks can be grouped into two main types: targeted and non-targeted. Targeted attacks aim to force the model to misclassify a specific instance (e.g., a particular node), whereas non-targeted attacks aim to degrade the overall performance of the model across the entire graph dataset. In [2], the authors are among the first to cast targeted attacks as a combinatorial optimization problem and to employ reinforcement learning to determine edge insertions and deletions that cause misclassification of a chosen node. In [11], Nettack is introduced as a targeted attack that crafts adversarial graphs by maximizing the training loss of GNNs. Building on this line of work, Metattack is proposed in [3] as a non-targeted method that treats the graph structure as a hyperparameter and exploits meta-gradients to generate structural perturbations.
However, the weaknesses of such attacks are also apparent. It is observed in [4] that adversarial attacks mainly affect high-rank graph properties; accordingly, a low-rank approximation of the graph is first constructed by performing truncated singular value decomposition (TSVD) on the adjacency matrix, and this denoised graph is then exploited to train a more robust GNN model. Building on the idea of leveraging low-rank structure, Pro-GNN is further proposed in [5]. It jointly learns a cleaned graph and a robust GNN model under low-rank constraints imposed on the graph structure.
To address the weaknesses of existing attack methods, in this paper we propose a new high-strength yet stealthy attack framework. Specifically, we first perform spectral adversarial robustness evaluation on the graph structure to quantitatively estimate the vulnerability of each edge. This enables the attacker to precisely identify and target the weakest links, thereby achieving effective attacks by slightly perturbing the weights of these vulnerable edges, without adding, deleting, or rewiring any edges. We conduct extensive experiments on widely used classical GNN models with real-world datasets, which demonstrate the effectiveness of the proposed method. The key contributions of this work are summarized as follows:
  1. 1.
    To the best of our knowledge, this is the first algorithm that leverages spectral adversarial robustness evaluation to quantitatively assess the attack-resistance of each edge in a graph for selecting attack targets accordingly.
  2. 2.
    In contrast to existing attack methods that require altering the graph topology—such as inserting, deleting, or rewiring a small number of edges—our method enables effective attacks without any changes to edge connectivity. Specifically, it achieves effective attacks by only slightly adjusting the weights of a small portion of edges, thereby making the attack harder to detect.
@@ 3 Method
As shown in Figure 1, our method consists of four phases, which are detailed in the rest of this section.
Figure 1: Overview of the proposed method.
@@ 4 Experiments
In this section, we validate our framework on three of the most popular GNN models.
4.1 Experiment Setup
4.1.1 Models
We evaluate three representative graph neural network (GNN) architectures, as described below.
  • Graph Convolutional Network (GCN): A spectral-inspired message-passing model that updates node representations by aggregating and transforming features from local neighborhoods through stacked graph convolution layers.
  • Graph Attention Network (GAT): A message-passing model that learns attention coefficients to adaptively weight contributions from neighbors during aggregation.
  • GraphSAGE: An inductive neighborhood-aggregation model. In our implementation, we use a weighted mean aggregator so that messages from neighbors are averaged according to edge weights.
4.1.2 Implementation and Parameter Settings
  • GCN: We adopt a two-layer Graph Convolutional Network (GCN). The first GCNConv maps node features to a 64-dimensional hidden representation followed by a ReLU activation, and the second GCNConv maps hidden features to the number of classes.
  • GAT: We use a two-layer Graph Attention Network (GAT) with weighted attention. The first layer uses 4 heads with 8 hidden units per head (concatenated), followed by an ELU activation; the second layer uses a single head to output class logits. We set dropout to 0.2.
  • GraphSAGE: We adopt a two-layer GraphSAGE with a weighted mean aggregator. In each layer, neighbor features are aggregated by a weight-aware average so that edges with larger weights contribute more to the message passing. The first layer uses a 32-dimensional hidden size with BatchNorm and ReLU (dropout $p=0$), and the second layer outputs class logits.
We use CrossEntropyLoss for multi-class classification and optimize with Adam for 20 epochs. Unless otherwise stated, the learning rate is set to 0.005 and weight decay to $\mathbf{5\times 10^{-4}}$.
For all three GNN models, we set $k=50$ for the kNN graph used to capture the manifold of the latent representations, and we set the number of eigenvectors used to perform spectral embedding to 20.
For each model, we attack only $10\%$ of the edges in the graph.
4.1.3 Data Set
We evaluate the proposed method on Cora, a widely used citation-network benchmark in which nodes represent scientific publications and edges indicate citation relationships. Each node is described by a sparse bag-of-words feature vector and assigned to one of seven classes. The graph contains 2,708 nodes, 5,429 edges, and 1,433 features.
4.1.4 Attack Strategy
To make the attack imperceptible, we avoid structure-altering operations such as adding, deleting, or rewiring edges. Instead, we keep the graph topology unchanged: once the top $10\%$ of the most vulnerable edges have been identified by our Spectral Edge Attack (SEA) method, we perturb the graph only by slightly reweighting these edges. Since the Cora graph is treated as unweighted in our implementation (all edges initially have unit weight), we simply increase the weight of each selected (attacked) edge from 1 to 2, while keeping all other edge weights unchanged.
To validate the effectiveness of the SEA method, we construct a baseline by uniformly sampling the same number of edges from the graph and applying the same reweighting scheme. We then compare the performance degradation caused by SEA-selected edges with that caused by these randomly selected edges. The random sampling procedure for the baseline is repeated for 10 trials, and we report the mean degradation.
4.1.5 Experiment Result
The experimental results are summarized in Table 1. The column “Original” reports the classification accuracy (%) of each GNN model on the clean (unattacked) graph. The last two columns, $\downarrow$ SEA” and $\downarrow$ Random”, show the corresponding change in accuracy (in percentage points) after applying our SEA attack and the baseline random reweighting strategy, respectively, where positive values indicate accuracy degradation and negative values indicate accuracy gain.
From Table 1, we observe that, across all three GNN models, SEA consistently leads to a more pronounced drop in accuracy than the baseline attack under the same reweighting budget. For GCN and GAT, SEA causes a degradation of 1.23 and 1.04 percentage points, respectively, whereas the random baseline has almost no adverse effect and even slightly improves the performance of GCN. This can be attributed to the fact that randomly increasing the weights of a small subset of edges does not systematically target critical decision-making regions of the graph and may occasionally act as a benign perturbation or regularization, thereby yielding marginal accuracy gains. In contrast, SEA deliberately focuses on spectrally identified weak links that are most sensitive to perturbations, which results in a consistent deterioration of model performance.
For GraphSAGE, SEA still induces a larger degradation than the random sampling strategy, but the absolute drop is smaller than that observed for GCN and GAT, and the random baseline has a comparatively larger impact on GraphSAGE than on the other two GNN models.This pattern may be explained by the weighted mean aggregation used by GraphSAGE. On the one hand, averaging over a larger neighborhood makes the model inherently less sensitive to any single edge, which limits the maximum impact that even spectrally selected edges can have. On the other hand, because edge weights directly scale neighbor contributions, randomly increasing a few weights can distort the local averaging more than in GCN or GAT, where degree normalization or learned attention tends to dampen such noise. Even in this setting, however, SEA still causes almost three times the degradation of the random baseline, highlighting its superiority in identifying truly critical edges for attacking GraphSAGE.
Figure 2 provides a visual comparison of the degradation caused by SEA and by the baseline attack for all three models. The figure clearly shows that SEA consistently yields stronger attack effects than random reweighting. This confirms that spectral adversarial robustness evaluation is able to pinpoint structurally weak links that are particularly vulnerable to such subtle perturbations.
   Method    Original    $\downarrow$ SEA    $\downarrow$ Random
   GCN    89.54    1.23    -0.25
   GAT    88.44    1.04    0.04
   GraphSAGE    89.17    0.62    0.23
Table 1: Classification accuracy (%) results.
Figure 2: Attack Effect Comparison.
@@ 5 Conclusion
In this paper, we proposed SEA (Spectral Edge Attack), a novel framework for attacking graph neural networks. Instead of altering the graph topology via edge insertion, deletion, or rewiring, SEA quantitatively assesses the vulnerability of each edge via spectral analysis and performs attacks by slightly reweighting only the most fragile edges. This allows SEA to degrade model performance while keeping the connectivity pattern unchanged, making the attack more difficult to detect. Experiments on three representative GNN architectures (GCN, GAT, and GraphSAGE) show that, under the same perturbation budget, SEA consistently causes larger accuracy degradation than randomly reweighting edges. These results demonstrate that SEA is effective at pinpointing structurally weak links and provides a principled basis for designing subtle, unnoticeable attacks on GNNs.